1. Meir, Benson, and Lau are partners and share income and loss in a 2:3:5 ratio...

1.

Meir, Benson, and Lau are partners and share income and loss in a 2:3:5 ratio (in percents: Meir, 20%; Benson, 30%; and Lau, 50%). The partnership's capital balances are as follows: Meir, $78,000; Benson, $119,000; and Lau, $203,000. Benson decides to withdraw from the partnership.

1. Prepare the journal entry to record Benson's withdrawal under each independent assumptions. (Do not round intermediate calculations.)

(a) Benson sells her interest to North for $160,000 after North is approved as a partner; (b) Benson gives her interest to a son-in-law, Schmidt, and Schmidt is approved as a partner; (c) Benson is paid $119,000 in partnership cash for her equity; (d) Benson is paid $157,000 in partnership cash for her equity; and (e) Benson is paid $15,500 in partnership cash plus equipment recorded on the partnership books at $35,500 less its accumulated depreciation of $11,600.

2. 2. Assume that Benson does not retire from the partnership described in Part 1. Instead, Rhode is admitted to the partnership on February 1 with a 25% equity. Prepare journal entries to record Rhode’s entry into the partnership under each separate assumption: Rhode invests (a) $133,333; (b) $97,333; and (c) $174,666. (Do not round your intermediate calculations.)
Answer #1

1.

a.

Benson Capital a/c Dr 119000
To North Capital a/c 119000

(Benson received $ 160,000 from North which is outside the firm.

Hence no entry is books

b.

Benson capital a/c Dr 119000
To Schmidt capital a/c 119000
(Transfer of capital)

c.

Benson capital a/c Dr 119000
To cash e/c 119000
(capital balance paid off and settled)

d.

Benson capital a/c Dr 119000
Meir Capital a/c Dr (38000*2/(2+5)) 10857
Lau Capital a/c Dr (38000*5/(2+5)) 27143
To Cash a/c 157000
(capital balance paid off along with goodwill of 38000 (ie 157000-119000)

e.

Benson Capital a/c Dr 119000
Accumulated depreciation Dr 11600
To Cash a/c 15500
To Equipment 35500

To Meir Capital a/c \

(119000+(35500-11600+15500))*2/7

22743
To Lau Capital (119000-(35500-11600+15500)*5/7 56857
(being equipment and cash paid off on settlement and balance distributed to existing partners

2. Rhode is admitted for 25%. So remaining share of exisiting partners is 75%. So Rhodes equivalent capital => (78000+119000+203000)*25/75 = 133,333.

a.

Cash a/c Dr 133333
To Rhodes Capital a/c 133,333
(Being partner admitted with the capital contribution received)

b.

Cash a/c Dr 97333
Meir Capital a/c Dr (133,333-97333 = 36000*2/10) 7200
Benson Capital a/c Dr (36000*3/10) 10800
Lau capital a/c Dr (36000*5/10) 18000
To Rhodes Capital a/c 133,333
(Being partner admitted with the capital contribution received and bonus trfd from exisiting partners' capital a/c's)

c.

Cash a/c Dr 174666
To Meir Capital a/c Dr (133,333-97333 = 36000*2/10) 8267
Benson Capital a/c Dr (36000*3/10) 12400
Lau capital a/c Dr (36000*5/10) 20666
To Rhodes Capital a/c 133,333
(Being partner admitted with the capital contribution received and bonus trfd to exisiting partners' capital a/c's)